# The Road to Trust: Building Enclaves within Confidential VMs

### v2

@@
Abstract.
Integrity is critical for maintaining system security, as it ensures that only genuine software is loaded onto a machine. Although confidential virtual machines (CVMs) function within isolated environments separate from the host, it is important to recognize that users still encounter challenges in maintaining control over the integrity of the code running within the trusted execution environments (TEEs). The presence of a sophisticated operating system (OS) raises the possibility of dynamically creating and executing any code, making user applications within TEEs vulnerable to interference or tampering if the guest OS is compromised.
To address this issue, this paper introduces NestedSGX, a framework which leverages virtual machine privilege level (VMPL), a recent hardware feature available on AMD SEV-SNP to enable the creation of hardware enclaves within the guest VM. Similar to Intel SGX, NestedSGX considers the guest OS untrusted for loading potentially malicious code. It ensures that only trusted and measured code executed within the enclave can be remotely attested. To seamlessly protect existing applications, NestedSGX aims for compatibility with Intel SGX by simulating SGX leaf functions. We have also ported the SGX SDK and the Occlum library OS to NestedSGX, enabling the use of existing SGX toolchains and applications in the system. Performance evaluations show that context switches in NestedSGX take about 32,000 – 34,000 cycles, approximately 1.33×1.$3$1.54×1.54$\times$1.54 × higher than that of Intel SGX. NestedSGX incurs minimal overhead in most real-world applications, with an average overhead below 2% for computation and memory intensive workloads and below 15.68% for I/O intensive workloads.
copyright: noneconference: ACM Conference on Computer and Communications Security; Due 15 May 2019; London, TBDjournalyear: 2019
1. Introduction
@@
At the core of the confidential computing stack is the trusted execution environment (TEE), which isolates the code and data of a confidential workload from other code running on the system, even at the highest privilege levels. In particular, Intel SGX (sgx, 2022) provides protection for the trusted components of an application, known as enclaves, allowing only measured code to run within the TEEs [^1] Intel SGXv2 (McKeen et al., 2016) enables dynamic changes to the page attribute and dynamic loading of code, but with strict controls in place.. On the other hand, VM-based TEEs, such as AMD SEV (Sev-Snp, 2020), Intel TDX (Intel, 2020b), and ARM CCA (arm, 2022) offer a fully backward-compatible confidential virtual machine (CVM), enabling the execution of existing applications without modifications. However, placing entire VMs within TEEs is generally considered less desirable: a VM image is far more than just a kernel and an application – it includes a large number of system services. Consequently, it remains uncertain whether it is more secure than running the software on premises or on existing cloud infrastructure. Therefore, an interesting question still remains open for VM-based TEEs:
@@
How can users attest and establish trust in applications running within VM-based TEEs, given the dynamic nature of building, loading, modification, and execution of arbitrary code within the CVM, if the guest OS might be compromised?
@@
A seminal work called vSGX (Zhao et al., 2022) was developed to virtualize Intel SGX on AMD SEV, ensuring full binary compatibility and enabling the execution of unmodified SGX programs in CVMs. vSGX adopts a two-VM approach where one VM handles the untrusted application while the other hosts the enclave. However, this emulation of SGX leaf instructions incurs significant overhead, particularly for control flow transfers across the boundary of the two VMs (e.g., EENTER and EEXIT). For instance, an empty ECall operation on vSGX takes approximately 1.5ms, which is around 160×160$\times$160 × slower than on native Intel SGX. This slow context switch severely limits the service throughput achievable within the enclave. In practical scenarios, this limitation becomes a critical performance bottleneck. To illustrate, vSGX is 6×6$\times$6 × slower than the baseline on the cURL benchmark, and launching a 256 MB size vSGX enclave takes about 5 minutes.
@@
Figure 1. Comparison between NestedSGX and other TEEs. Firstly, NestedSGX offers a layered protection mechanism against both the host VMM and the guest OS within the CVM. Secondly, NestedSGX facilitates the capability to measure and attest the integrity of enclave code in environments where the guest OS is not fully trusted.
@@
Design. We take a further step in this direction by introducing our system, called NestedSGX. NestedSGX follows a defense-in-depth approach to attesting applications (as depicted in Figure 1), where both the enclaves and the guest OS run within the same CVM. Specifically, the trusted portion of an application (i.e., the enclave) is loaded and executed within an isolated environment, separated from the feature-rich OS, similar to the enclave abstraction of Intel SGX. During the loading process, the trusted portion is measured and can be attested to a remote user. The core root of trust for measurement (CRTM) and root of trust for reporting (RTR) rely on a small segment of trusted code, known as the security monitor. This security monitor operates at the highest privilege level within the CVM, and its integrity is also measured and attested with the assistance of hardware-based TEE. This approach empowers us to dynamically measure and attest the application code, enabling us to bypass the complexities associated with the guest OS.
@@
In more details, our approach leverages the privilege separation offered by AMD SEV-SNP, particularly the virtual machine privilege level (VMPL), to establish isolated guest physical space exclusively occupied by the security monitor and the enclaves. With VMPL, the vCPUs within the CVM are configured to operate at different privilege levels. Each privilege level is associated with specific access permissions to the guest physical memory pages. This capability enables us to assign distinct privilege levels to different software components running within the CVM. Particularly, the security monitor and the enclave operate at the highest VMPL (VMPL0), in the kernel mode and user mode respectively. On the other hand, the guest OS and the untrusted part of the application (App) are required to run in the kernel mode and user mode at a lower VMPL (e.g., VMPL1). This design ensures that the guest OS does not have access to the memory allocated for VMPL0.
@@
To facilitate the protection of existing applications, NestedSGX prioritizes compatibility with Intel SGX. This involves simulating SGX leaf functions to manage the entire life cycle of enclaves, including creation, memory management, teardown, and context switches. Furthermore, we have successfully ported the Intel SGX SDK (Intel, 2020a), the Rust SGX SDK (Wang et al., 2019) as well as the Occlum library OS (Shen et al., 2020) to the NestedSGX framework, allowing for smooth integration of existing SGX toolchains and applications. In addition, we have implemented the HotCalls (Weisse et al., 2017) optimization in Occlum, which allows specific OCALLs to be handled asynchronously within the enclave without exiting it. This optimization has led to improved performance for applications running on Occlum without requiring any modifications to their source code. Unlike vSGX, we stress that full binary compatibility with SGX is a non-goal of our paper.
@@
We have implemented NestedSGX on top of AMD’s Linux Secure VM Service Module (SVSM) framework (SVSM, 2022; AMD, 2023c), without necessitating any modifications of the host software, such as KVM. We conducted a series of benchmarking tests on commercial AMD SEV-SNP hardware and observed that the cost of world switches is about 1.33×1.$3$1.54×1.54$\times$1.54 × higher than Intel SGX, which is nearly two orders of magnitude faster than vSGX. Moreover, NestedSGX introduces negligible overhead in most real-world applications, with less than 2% overhead for memory and computation-intensive tasks and less than 15.68% overhead for I/O-intensive tasks. Generally, NestedSGX exhibits moderate costs that are comparable to other enclave systems, such as Intel SGX. Consequently, we view it as a promising solution for addressing the critical challenge of performing secure computations within CVMs.
@@
    Design. We present the design of NestedSGX, incorporating the VMPL feature within AMD SEV-SNP. NestedSGX addresses the challenge of establishing trust in the applications while minimizing reliance on the potentially compromised guest OS.
@@
    Implementations and evaluations. We implemented NestedSGX on commercial hardware. The evaluation results demonstrate that the overhead incurred by NestedSGX is comparable to that of Intel SGX, affirming its efficiency and viability.
2. Background
2.1. Intel SGX
@@ 2.2. SEV-SNP and VMPL
SEV-SNP, introduced in 2020 (Sev-Snp, 2020), is the third generation of SEV and provides enhanced security against malicious manipulation of page mappings by the host. One of its key features is the Reverse Map Table (RMP), a structure located in secure memory that maps system physical addresses (sPAs) to guest physical addresses (gPAs). The RMP serves as a metadata table managed by the AMD-SP and plays a critical role in tracking the ownership of each system physical page to prevent the host from writing to encrypted guest pages. It establishes a global one-to-one mapping between sPAs and gPAs, ensuring that a page cannot be simultaneously mapped into multiple guests or multiple times within a single guest.
Virtual Machine Privilege Level (VMPL) is an optional feature within the SEV-SNP architecture that allows a guest VM to divide its address space into four distinct levels. These levels serve as hardware-isolated abstraction layers for the CVM, with VMPL0 representing the highest privilege level and VMPL3 representing the lowest. Each hardware context, known as a Virtual Machine Save Area (VMSA), is associated with a specific VMPL. Moreover, different memory pages assigned to a guest can have varying permissions based on the VMPL. VMPLs are utilized to implement additional page permission checks and are independent of other x86 security features.
The RMP maintains records of the read, write, and execution permissions for each guest physical page across different VMPLs. During the translation of a virtual address to a physical address by the CPU or IOMMU, an RMP check is typically conducted to determine the relevant permissions and ownership of each physical memory page. The guest has the ability to modify VMPL permissions using the RMPADJUST instruction. This instruction allows a higher privilege VMPL to adjust the permissions of a lower privilege VMPL. It grants the guest the flexibility to modify and manage the permissions of different VMPLs based on its specific requirements. In addition to SEV-SNP, AMD provides the Secure Virtual Machine Service Module (SVSM) framework (SVSM, 2022), which is a small piece of code running at the highest privilege level (VMPL0) to provide security services (such as virtual TPM) to the rest of the guest.
@@
2.3. Threat model
NestedSGX is built on top of SEV-SNP and follows a similar threat model, which considers the host software as a potential source of harm or compromise. We place our trust in the underlying SEV-SNP hardware to protect the CVM from direct observation or tampering by the host software or other CVMs. Specifically, we do not consider the recent attacks on SEV-SNP, such as the CacheWrap attack (Zhang et al., 2024), the WeSee attack (Schlüter et al., 2024a) and the Heckler attack (Schlüter et al., 2024b).
@@ 2.3. Threat model
Furthermore, this paper does not cover scenarios in which the enclave intentionally reveals secrets or the security risks associated with unsafe code, such as memory safety vulnerabilities within the enclave or attacks originating from the interface between the enclave and the application (e.g., COIN (Khandaker et al., 2020) or IAGO (Checkoway and Shacham, 2013) attacks). However, it is important to highlight that NestedSGX ensures that the enclave cannot compromise the operations or integrity of the guest OS, even in the presence of vulnerabilities. The security monitor operates in kernel mode at VMPL0 and is responsible for enforcing memory isolation between mutually distrustful components. Additionally, the security monitor serves as the core root of trust for measurement (CRTM). To establish a foundation of trust, it is assumed that the security monitor is fully trusted.
@@
This paper does not cover side channel attacks, including traditional cache side channel attacks (Liu et al., 2015), page table based attacks (Xu et al., 2015; Wang et al., 2017; Van Bulck et al., 2017) or transient execution attacks (Kocher et al., 2020; Van Bulck et al., 2018; Van Schaik et al., 2019). We consider protection against these types of attacks as orthogonal to our design. Furthermore, since the host and the guest OS are free to not schedule the enclave for execution, Denial-of-Service (DoS) attacks are also out of scope.
3. Design
3.1. Overview
@@
Figure 2. NestedSGX overview. The security monitor and the enclave operate in the kernel and user mode at VMPL0, while the guest OS and App operate in the kernel and user mode at VMPL1. The transitions between VMPL0 and VMPL1 occur via the untrusted host.
@@
As shown in Figure 2, NestedSGX encompasses the following components within the CVM. The application is divided into the trusted part (enclave) and the untrusted part (App). The security monitor is the root of trust for measurement and reporting of the enclave. It is also responsible for enforcing security isolation within NestedSGX. The guest OS (such as popular Linux distributions) offers essential features to applications, such as file systems, networking, device drivers, and language runtimes. To let the users maintain control over the code running inside the CVM and bootstrap trust to the enclave within a feature-rich guest OS, NestedSGX is designed to fulfill the following requirements.
@@
Firstly, NestedSGX ensures that any transitions between the trusted and untrusted components are mediated by the security monitor. This guarantees that sensitive information is properly sanitized and safeguarded. To leverage existing SGX toolchains, NestedSGX employs the SGX programming model to manage the enclave’s life cycle. The details for secure enclave state transitions will be presented in Sec. 3.2.
@@
Secondly, the system establishes isolated memory regions exclusively used by the security monitor, the enclaves and the guest OS respectively. This ensures that the enclave and the security monitor are secure even in scenarios where the guest OS operates in kernel mode. Specifically, the entire gPA space is divided to two parts: a continuous and fixed region exclusively used by VMPL0, and the remaining portion used by VMPL1. The security monitor and the enclave are executed at VMPL0, while the guest OS and App operate at VMPL1. Additionally, our design guarantees that the enclave cannot access the memory of the guest OS, despite running at a higher VMPL. The details of the memory isolation scheme will be presented in Sec. 3.3.
@@
Lastly, in NestedSGX, only authenticated code and data can be loaded and processed within the enclave. To achieve this, NestedSGX enforces a mechanism where the enclave can only be loaded by the trusted security monitor, which also performs the measurement of the enclave’s integrity. While the guest OS retains the freedom to load and execute any code within its own address space, these code segments are not affirmed by the security monitor and therefore will not be attested. The details of enclave measurement and attestation will be presented in Sec. 3.4.
Table 1. Comparison with competing approaches.
@@
vSGX (Zhao et al., 2022)
Veil (Ahmad et al., 2024)
@@
Comparison with competing approaches (Table 1). vSGX (Zhao et al., 2022) effectively virtualizes Intel SGX on AMD SEV, ensuring complete binary compatibility and the execution of unmodified SGX programs. vSGX employs two separated CVMs to accommodate the untrusted application and the enclave. In contrast, NestedSGX adopts a more intuitive design that utilizes different VMPLs for memory isolation. The NestedSGX design offers the following benefits.
@@
Secondly, with vSGX, only one enclave can be hosted within the EVM, which restricts the maximum number of concurrently-running enclaves on the hardware platform. This limitation arises because SEV associates ASIDs with VMs’ memory encryption keys, and the ASID bit is limited (4 bits (Gu et al., 2020)). In contrast, NestedSGX has the ability to support an unlimited number of enclaves.
More importantly, NestedSGX does not serve as a direct substitute for vSGX. On vSGX, the enclave thread and application thread operate with different vCPUs in separated VMs. Specifically, the application thread at the AVM is paused on EENTER; upon receiving the EENTER request, the EVM creates a thread to handle it in the EVM. On NestedSGX, the enclave thread and application thread operate with the same vCPU, while the vCPU contexts associated with different VMPLs are saved in different VMSAs. Consequently, managing the entry and exit of enclaves on NestedSGX necessitates meticulous handling of vCPU state transitions. As an evidence, AEX is relatively straightforward on vSGX: vSGX simply puts the enclave thread to sleep on AEX, then wakes it and resumes the execution directly on ERESUME. On NestedSGX the enclave context needs to be saved on AEX and recovered on ERESUME.
Concurrent to our work, Ahmad et al. propose Veil (Ahmad et al., 2024) (to be presented at ASPLOS 2024) as a service framework running at VMPL0. The concept of Veil is quite similar to AMD’s Secure VM Service Module (SVSM). As one of the use cases, Veil offers protection to the entire applications within a separated VMPL, facilitating redirection of system calls and interrupts. However, unlike our approach, it does not focus on protecting a trusted portion of the application or strive for compatibility with Intel SGX. Currently, Veil lacks support for multi-threading and exception handling. In contrast, NestedSGX leverages an existing SGX toolchain by emulating most SGX instructions, thereby naturally providing support for multi-threading and exception handling.
3.2. Enclave life cycle management

### v3

@@
Abstract
Integrity is critical for maintaining system security, as it ensures that only genuine software is loaded onto a machine. Although confidential virtual machines (CVMs) function within isolated environments separate from the host, it is important to recognize that users still encounter challenges in maintaining control over the integrity of the code running within the trusted execution environments (TEEs). The presence of a sophisticated operating system (OS) raises the possibility of dynamically creating and executing any code, making user applications within TEEs vulnerable to interference or tampering if the guest OS is compromised. To address this issue, this paper introduces NestedSGX, a framework which leverages virtual machine privilege level (VMPL), a recent hardware feature available on AMD SEV-SNP to enable the creation of hardware enclaves within the guest VM. Similar to Intel SGX, NestedSGX considers the guest OS untrusted for loading potentially malicious code. It ensures that only trusted and measured code executed within the enclave can be remotely attested. To seamlessly protect existing applications, NestedSGX aims for compatibility with Intel SGX by simulating SGX leaf functions. We have also ported the SGX SDK and the Occlum library OS to NestedSGX, enabling the use of existing SGX toolchains and applications in the system. Performance evaluations show that context switches in NestedSGX take about 32,000 – 34,000 cycles, approximately 1.9×1.9$\times$1.9 ×2.1×2.1$\times$2.1 × higher than that of Intel SGX. NestedSGX incurs minimal overhead in most real-world applications, with an average overhead below 2% for computation and memory intensive workloads and below 15.68% for I/O intensive workloads.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 24 February - 28 February 2025, San Diego, CA, USA ISBN x-xxxxxx-xx-x https://dx.doi.org/xx.xxxxx/ndss.2025.xxxxx www.ndss-symposium.org
I Introduction
@@
At the core of the confidential computing stack is the trusted execution environment (TEE), which isolates the code and data of a confidential workload from other code running on the system, even at the highest privilege levels. In particular, Intel SGX [5] provides protection for the trusted components of an application, known as enclaves, allowing only measured code to run within the TEEs [^1] Intel SGXv2 [39] enables dynamic changes to the page attribute and dynamic loading of code, but with strict controls in place.. On the other hand, VM-based TEEs, such as AMD SEV [47], Intel TDX [30], and ARM CCA [4] offer a fully backward-compatible confidential virtual machine (CVM), enabling the execution of existing applications without modifications. However, placing entire VMs within TEEs is generally considered less desirable: a VM image is far more than just a kernel and an application – it includes a large number of system services. It remains uncertain whether it is more secure than running the software on premises or on existing cloud infrastructure. Therefore, an interesting question still remains open for VM-based TEEs:
@@
How can users attest and establish trust in applications running within a CVM, given the dynamic nature of building, loading, modification, and execution of arbitrary code within the CVM, if the guest OS might be compromised?
@@
A seminal work called vSGX [62] was developed to virtualize Intel SGX on AMD SEV, ensuring full binary compatibility and enabling the execution of unmodified SGX programs in CVMs. vSGX adopts a two-VM approach where one VM handles the untrusted application while the other hosts the enclave. However, this emulation of SGX leaf instructions incurs significant overhead, particularly for control flow transfers across the boundary of the two VMs (e.g., EENTER and EEXIT). For instance, an empty ECall operation on vSGX takes approximately 1.5ms, which is around 160×160$\times$160 × slower than on native Intel SGX. This slow context switch severely limits the service throughput achievable within the enclave. In practical scenarios, this limitation becomes a critical performance bottleneck. To illustrate, vSGX is 6×6$\times$6 × slower than the baseline on the cURL benchmark, and launching a 256 MB size vSGX enclave takes about 5 minutes.
@@
Figure 1: Comparison between NestedSGX and other TEEs. NestedSGX offers a layered protection mechanism against both the host VMM and the guest OS within the CVM.
@@
Design. We take a further step in this direction by introducing our system, called NestedSGX. NestedSGX follows a defense-in-depth approach to attesting applications (as depicted in Fig. 1), where both the enclaves and the guest OS run within the same CVM. Specifically, the trusted portion of an application (i.e., the enclave) is loaded and executed within an isolated environment, separated from the feature-rich OS, similar to the enclave abstraction of Intel SGX. During the loading process, the trusted portion is measured and can be attested to a remote user. The core root of trust for measurement (CRTM) and root of trust for reporting (RTR) rely on a small segment of trusted code, known as the security monitor. This security monitor enforces memory isolation within the CVM, and its integrity is also measured and attested with the assistance of hardware-based TEE. This approach empowers us to dynamically measure and attest the application code, enabling us to bypass the complexities associated with the guest OS. To facilitate the protection of existing applications, NestedSGX prioritizes compatibility with Intel SGX. This involves simulating SGX leaf functions with a trusted SGX emulation layer to manage the entire life cycle of enclaves, including creation, memory management, teardown, and context switches.
@@
In more details, our approach leverages the privilege separation offered by AMD SEV-SNP, particularly the virtual machine privilege level (VMPL), to establish isolated guest physical space exclusively occupied by the security monitor and the enclaves. With VMPL, the vCPUs within the CVM are configured to operate at different privilege levels. Each privilege level is associated with specific access permissions to the guest physical memory pages. This capability enables us to assign distinct privilege levels to different software components running within the CVM. Particularly, the security monitor, the SGX emulation layer and the enclave operate at the higher VMPL (e.g., VMPL0), in the kernel mode and user mode respectively. On the other hand, the guest OS and the untrusted part of the application (App) are required to run in the kernel mode and user mode at a lower VMPL (e.g., VMPL1). This design ensures that the guest OS does not have access to the memory allocated for VMPL0.
@@
We implemented NestedSGX on top of AMD’s Linux Secure VM Service Module (SVSM) framework [16], without necessitating any modifications of the host software, such as KVM. Furthermore, we have successfully ported the Intel SGX SDK [29], the Rust SGX SDK [56] as well as the Occlum library OS [48] to the NestedSGX framework, allowing for smooth integration of existing SGX toolchains and applications. In addition, we have implemented the HotCalls [58] optimization in Occlum, which allows specific OCALLs to be handled asynchronously within the enclave without exiting it. This optimization has led to improved performance for applications running on Occlum without requiring any modifications to their source code. Unlike vSGX, we stress that full binary compatibility with SGX is a non-goal of our paper.
@@
We conducted a series of benchmarking tests on commercial AMD SEV-SNP hardware and observed that the cost of world switches is about 1.9×1.9$\times$1.9 ×2.1×2.1$\times$2.1 × higher than Intel SGX, which is nearly two orders of magnitude faster than vSGX. Moreover, NestedSGX introduces negligible overhead in most real-world applications, with less than 2% overhead for memory and computation-intensive tasks and less than 15.68% overhead for I/O-intensive tasks. Generally, NestedSGX exhibits moderate costs that are comparable to other enclave systems, such as Intel SGX. Consequently, we view it as a promising solution for addressing the critical challenge of performing secure computations within CVMs.
@@
    Design. We present the design of NestedSGX, incorporating the VMPL feature within AMD SEV-SNP. NestedSGX addresses the challenge of establishing trust for applications while minimizing reliance on the potentially compromised guest OS.
@@
    Implementations and evaluations. We implemented NestedSGX on commercial hardware. The evaluation demonstrates the overhead incurred by NestedSGX is comparable to that of Intel SGX, affirming its efficiency and viability.
Roadmap. The rest of the paper is organized as follows. We provide necessary background information about Intel SGX and SEV-SNP, along with the threat model in Sec. II. We present the detailed design of NestedSGX in Sec. III, followed by implementation details in Sec. IV. The security analysis and performance evaluation of NestedSGX are provided in Sec. V and Sec. VI, respectively. We discuss possible extensions and related works of NestedSGX in Sec. VII and Sec. VIII. Sec. IX concludes the paper.
II Background
II-A Intel SGX
@@ II-B SEV-SNP and VMPL
SEV-SNP, introduced in 2020 [47], is the third generation of SEV and provides enhanced security against malicious manipulation of page mappings by the host. One of its key features is the Reverse Map Table (RMP), a structure located in secure memory that maps system physical addresses (sPAs) to guest physical addresses (gPAs). The RMP serves as a metadata table managed by the AMD-SP and plays a critical role in tracking the ownership of each system physical page to prevent the host from writing to encrypted guest pages. It establishes a global one-to-one mapping between sPAs and gPAs, ensuring that a page cannot be simultaneously mapped into multiple guests or multiple times within a single guest.
Virtual Machine Privilege Level (VMPL) is an optional feature within the SEV-SNP architecture that allows a CVM to divide its address space into four distinct levels. These levels serve as hardware-isolated abstraction layers for the CVM, with VMPL0 representing the highest privilege level and VMPL3 representing the lowest. Each hardware context, known as a Virtual Machine Save Area (VMSA), is associated with a specific VMPL. Moreover, different memory pages assigned to a guest can have varying permissions based on the VMPL. VMPLs are utilized for additional page permission checks and are independent of other x86 security features.
The RMP maintains records of the read, write, and execution permissions for each guest physical page across different VMPLs. During the translation of a virtual address to a physical address by the CPU or IOMMU, an RMP check is typically conducted to determine the relevant permissions and ownership of each physical memory page. The guest has the ability to modify VMPL permissions using the RMPADJUST instruction. This instruction allows a higher privilege VMPL to adjust the permissions of a lower privilege VMPL. It grants the guest the flexibility to modify and manage the permissions of different VMPLs based on its specific requirements. In addition to SEV-SNP, AMD provides the Secure Virtual Machine Service Module (SVSM) framework [50], which is a small piece of code running at the highest privilege level (VMPL0) to provide security services (such as virtual TPM) to the rest of the guest.
@@
II-C Threat model
NestedSGX is built on top of SEV-SNP and follows a similar threat model, which considers the host software as a potential source of harm or compromise. We place our trust in the underlying SEV-SNP hardware to protect the CVM from direct observation or tampering by the host software or other CVMs. Specifically, we do not consider the recent attacks on SEV-SNP, such as the CacheWrap attack [60], the WeSee attack [44] and the Heckler attack [45].
@@ II-C Threat model
Furthermore, this paper does not cover the security risks associated with unsafe code, such as memory safety vulnerabilities within the enclave [19], enclave malwares [52, 46], or attacks originating from the interface between the enclave and the application (e.g., COIN [32] or IAGO [22] attacks). However, it is important to highlight that NestedSGX ensures that the enclave cannot compromise the operations or integrity of the guest OS, even in the presence of vulnerabilities. The security monitor is responsible for enforcing memory isolation between mutually distrustful components. The SGX emulation layer serves as the core root of trust for measurement (CRTM). To establish a foundation of trust, it is assumed that the security monitor and the SGX emulation layer are trusted.
@@
This paper does not cover side channel attacks, including traditional cache side channel attacks [35], page table based attacks [59, 57, 54] or transient execution attacks [33, 53, 55]. We consider protection against these types of attacks as orthogonal to our design. Furthermore, since the host and the guest OS are free to not schedule the enclave for execution, Denial-of-Service (DoS) attacks are also out of scope.
III Design
III-A Overview
@@
Figure 2: NestedSGX overview. The security monitor, the SGX emulation layer and the enclave operate in the kernel and user mode at VMPL0, while the guest OS and App operate in the kernel and user mode at VMPL1. The transitions between VMPL0 and VMPL1 occur via the untrusted host.
@@
As shown in Fig. 2, NestedSGX encompasses the following components within the CVM. The application is divided into the trusted part (enclave) and the untrusted part (App). The security monitor is responsible for enforcing security isolation within NestedSGX. The SGX emulation layer emulates SGX instructions, and serves as the root of trust (RoT) for measurement and reporting of the enclave. The guest OS offers essential features to applications, such as file systems, networking, device drivers, and language runtime.
To enforce memory isolation, the security monitor runs at the highest privileged VMPL (i.e., VMPL0). On the other hand, we cannot place the SGX emulation layer in a VMPL lower than that of the guest OS (e.g., guest OS in VMPL1 and SGX emulation layer in VMPL2). Doing so would grant the guest OS the ability to generate attestation reports on behalf of the SGX emulation layer, potentially compromising the trust chain (please refer to Sec. III-D for details). Within our threat model, we consider the SGX emulation layer to be trusted. For the sake of brevity, throughout the remainder of the paper, we regard the SGX emulation layer as an integral component of the security monitor, both running at VMPL0. while the guest OS and App operate at a lower privileged VMPL (i.e., VMPL1). Within this design, we also place the enclaves within the same VMPL as the SGX emulation layer, in the user mode and kernel mode respectively, since if they were placed at different VMPLs, a context switch between the enclave and App would necessitate 2 costly VMPL switches. In practice, however, if running both the SGX emulation layer and the enclaves in VMPL0 raises security concerns, we can adapt NestedSGX slightly as follows: the security monitor still runs in VMPL0; the SGX emulation layer and the enclaves run in the kernel and user mode at VMPL1, and the guest OS and the App run in the lowest VMPL (e.g., VMPL3).
To let the users maintain control over the code running inside the CVM and bootstrap trust to the enclave within a feature-rich guest OS, NestedSGX is designed to fulfill the following requirements.
@@
Firstly, NestedSGX ensures that any transitions between the trusted and untrusted components are mediated by the security monitor. This guarantees that sensitive information is properly sanitized and safeguarded. To leverage existing SGX toolchains, NestedSGX employs the SGX programming model to manage the enclave’s life cycle. The details for secure enclave state transitions will be presented in Sec. III-B.
@@
Secondly, the system establishes isolated memory regions exclusively used by the security monitor, the enclaves and the guest OS respectively. This ensures that the enclave and the security monitor are secure even in scenarios where the guest OS operates in kernel mode. Specifically, the entire gPA space is divided to two parts: a continuous and fixed region exclusively used by VMPL0, and the remaining portion used by VMPL1. Additionally, our design guarantees that the enclave cannot access the memory of the guest OS, despite running at a higher VMPL. The details of the memory isolation scheme will be presented in Sec. III-C.
@@
Lastly, in NestedSGX, only authenticated code and data can be loaded and processed within the enclave. To achieve this, NestedSGX enforces a mechanism where the enclave can only be loaded by the trusted security monitor, which also performs the measurement of the enclave’s integrity. While the guest OS retains the freedom to load and execute any code within its own address space, these code segments are not affirmed by the security monitor and therefore will not be attested. The details of enclave measurement and attestation will be presented in Sec. III-D.
Table I: Comparison with competing approaches.
@@
vSGX [62]
Veil [13]
@@
Comparison with competing approaches (Table I). vSGX [62] effectively virtualizes Intel SGX on AMD SEV, ensuring complete binary compatibility and the execution of unmodified SGX programs. vSGX employs two separated CVMs to accommodate the untrusted application and the enclave. In contrast, NestedSGX adopts a more intuitive design that utilizes different VMPLs for memory isolation. The NestedSGX design offers the following benefits.
@@
Secondly, with vSGX, only one enclave can be hosted within the EVM, which restricts the maximum number of concurrently-running enclaves on the hardware platform. This limitation arises because SEV associates ASIDs with VMs’ memory encryption keys, and the ASID bit is limited. In contrast, NestedSGX has the ability to support an unlimited number of enclaves.
Concurrent to our work, Ahmad et al. proposed Veil [13] as a service framework running at VMPL0. As one use case, Veil offers protection to the entire applications within a separated VMPL, facilitating redirection of system calls and interrupts. However, its design has some limitations: (1) Exception handling requires enclave context information and is not supported; (2) On context switches (e.g., syscall and interrupt), VMPL switches to the guest OS are triggered by the enclave, and don’t follow the standard GHCB protocols, necessitating changes to the hypervisor; (3) Veil does not support SGX ecosystems. Addressing these limitations requires an SGX emulation layer to emulate SGX instructions and fulfill certain OS functionalities, such as mediating all enclave-App context switches and forwarding exceptions to the guest OS, within the user-land enclave’s VMPL. Designing and implementing this layer is challenging because its state – including general-purpose and segment registers, stack pointers – changes during context switches when handling various requests (e.g., forwarding exceptions or emulating SGX instructions), so when the system re-enters the enclave again, its state cannot be fully restored due to the change of the state in the emulation layer that handles the context switch. As a result, the handling of subsequent requests can lead to failures, such as faults and stack exhaustion.
Note that this challenge cannot be resolved by vSGX or Veil. vSGX runs the App and enclave in different VMs and can simply pause or activate one of them for context switching, but also incurs significant overheads for cross-VM communications during the switch. Context switches in Veil rely on hardware, which automatically saves and restores the contexts of the enclave and the guest OS within their respective VMSAs. But this mechanism cannot handle the switches triggered by exceptions.
Our solution introduces a unique design that models the states of the SGX emulation layer using a finite state machine (FSM). After each context switch, the layer restores its original state under the FSM’s guidance. We achieve this by pairing asynchronous and synchronous enclave entry and exit requests. We save the state of the SGX emulation layer before the entry request (EENTER/AEX) and restore it after processing the exit request, through a carefully designed restoration strategy based upon the FSM (e.g., restoring gs before rsp to balance the stack). In this way, the enclave state can be guaranteed to be correct no matter how complicated the request sequence could become (e.g., arbitrary exceptions between EENTER and EEXIT). This design, along with our thorough implementation of exception handling and multithreading, differentiates NestedSGX from both Veil and vSGX.
III-B Enclave life cycle management